Based on a​ poll, 60​% of adults believe in reincarnation. Assume that 6 adults are randomly​ selected, and find the indicated probability. What is the probability that all of the selected adults believe in​ reincarnation?




1
Expert's answer
2021-12-07T11:20:26-0500

"p=0.6 \\

q = 1-p = 1 -0.6 = 0.4 \\

n= 6"

Using Binomial distribution

"P(X=x) = \binom{n}{x}p^xq^{n-x} \\

P(X=6) = \binom{6}{6} \times 0.6^6 \times 0.4^{6-6} \\

= 1 \times 0.6^6 \times 1 \\

= 0.046656"